OpenStudy (anonymous):

What is the slope of the line that contains the points (1,5) and (3,2)?

OpenStudy (anonymous):

The slope is -1.5

OpenStudy (anonymous):

no @*shadow*

OpenStudy (anonymous):

There's a 2 difference in the (x) coordinate 1 to 3. There's a 3 difference in the (x) coordinate 5 to 2. If you're looking to change 3 numbers in 2 numbers.. 3/2 = 1.5 and since the slope is negative = -1.5.. My logic.

OpenStudy (anonymous):

I'm pretty sure, I'm right.

OpenStudy (anonymous):

Not really the way to figure it out.. but it's my way :)

OpenStudy (lyrae):

y = mx + b slope = m \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

Point1: (3,2) Point2: (1,5) \[m=\frac{ y _{2}-y _{1} }{ x _{2}-_x{1}}\] \[m=\frac{ 2-5}{3-1 }\] \[m=-1.5\]

OpenStudy (anonymous):

@lozada796, the answer is -1.5...